OpenStudy (gorica):

Foucault's pendulum - why the last equation (Newton's second law) is equal to zero?

OpenStudy (dumbcow):

\[r = <x,y,z>\] \[ak = <0,0,a>\] \[r-ak = <x,y,z-a>\] \[|r-ak|^{2} = x^{2}+y^{2}+(z-a)^{2}\] \[=x^{2}+y^{2}+z^{2}+a^{2}-2az\] im doing something wrong

OpenStudy (dumbcow):

by definition \[|r|^{2} = a^{2}\] thus you are saying ak = 0 ?

OpenStudy (gorica):

 I wrote picture like this. ak can't be zero since a is length of a string.